# From Optimization to Control: Quasi Policy Iteration

### v3

@@ Abstract.
The authors are with Delft Center for Systems and Control, Delft University of Technology, Delft, The Netherlands. Email: {M.A.SharifiKolarijani, P.MohajerinEsfahani}@tudelft.nl.
This research is part of a project that has received funding from the European Research Council (ERC) under the grant TRUST-949796.
@@ 1. Introduction
The problem of control, or the decision-making problem as it is also known within the operations research community, has been the subject of much research since the introduction of the Bellman principle of optimality in the late 1950s [4]. In particular, the connection between control algorithms for Markov decision processes (MDPs) and optimization algorithms has been noticed since the late 1970s [43]: Value iteration (VI) [4] can be seen as an instance of gradient descent (GD) algorithm, and policy iteration (PI) [27] is an instance of the Newton method (NM).
@@
More recent works have used the relationship mentioned above to develop new control algorithms, with faster convergence and/or lower complexity, inspired by their counterparts for solving optimization problems [23, 52]. In case of model-based (a.k.a. planning, with access to the model of the MDP) algorithms, the combination of the VI algorithm with Polyak momentum [40], Nesterov acceleration [38], and Anderson acceleration [2] have been explored in [22] and [56]. More recently, Halpern anchoring acceleration [24] has been used to introduce the Anchored VI algorithm [33], which in particular exhibits an $\operatorname{\mathcal{O}}(1/k)$-rate for large values of discount factor and even for $\gamma=1$. Also of notice is the Generalized Second-Order VI algorithm [28] which applies NM on a smoothed version of the Bellman operator. For model-free (a.k.a. learning, with access to samples from the MDP) algorithms, Speedy Q-Learning [20], Momentum Q-Learning [54], and Nesterov Stochastic Approximation [14] are among the algorithms that use the idea of momentum for accelerating stochastic GD algorithm [55, 30, 34, 1] in order to achieve a better rate of convergence compared to standard Q-learning (QL). Moreover, the Zap Q-Learning algorithm [15] can be thought of as a second-order learning algorithm which was inspired by the stochastic Newton-Raphson (SNR) algorithm [45].
@@
We note that all of the aforementioned connections have been focused on finite state-action MDPs. When it comes to infinite (continuous) state-action spaces, except in special cases such as linear–quadratic regulators (LQR), one needs to resort to finite-dimensional approximation techniques for computational purposes. This approximation may be at the modeling level by aggregation (discretization) of the state and action spaces, which readily falls into the finite state-action MDP setting [6, 41]. Alternatively, one may directly approximate the value function via finite parametrization and minimizing (a proxy of) the residual of its fixed-point characterization based on the Bellman principle of optimality [7, 50]. Examples of such include linear parameterization [10, 51], or nonlinear parameterization with, for instance, neural network architectures [9, 46, 49] or max-plus approximation [5, 21, 32, 31, 36, 35]. We also note that there is an alternative characterization of the original function as the solution to an infinite-dimensional linear program [25], paving the way for approximation techniques via finite tractable convex optimization [13, 26, 37]. With this view of the literature, it is worth noting that one can cast almost all of these approximation techniques as the solution to a finite-dimensional fixed-point or convex optimization problem.
@@ 3. Quasi-Policy Iteration (QPI)
where $f:\mathbb{R}^{\ell}\rightarrow\mathbb{R}$ is twice continuously differentiable and strongly convex, and let $g(x)=\nabla f(x)$ and $H(x)=\nabla^{2}f(x)$ be the gradient and Hessian of $f$ evaluated at $x$. Then, by defining $g(v)\coloneqq v-T(v)$, one can see that the VI update rule [4]
@@
where $\alpha_{k}$ is the step-size. Similarly, by defining $H(v)\coloneqq I-\gamma P^{\pi_{v}}$, the PI update rule [42, Prop. 6.5.1]
@@
The VI algorithm converges linearly with rate $\gamma$ owing to the fact that the operator $T$ is a $\gamma$-contraction in the $\infty$-norm [42, Prop. 6.2.4]. The PI algorithm, on the other hand, outputs the optimal policy in a finite number of iterations [42, Thm. 6.4.2]. Moreover, the algorithm has a local quadratic rate of convergence when initiated in a small enough neighborhood around the optimal solution [8, 18]. The faster convergence of PI compared to VI, however, comes with a higher per-iteration computational complexity: The per-iteration complexities of VI and PI are $\operatorname{\mathcal{O}}(n^{2}m)$ and $\operatorname{\mathcal{O}}(n^{2}m+n^{3})$, respectively. The extra $\operatorname{\mathcal{O}}(n^{3})$ complexity is due to the policy evaluation step, i.e., solving a linear system of equations; see also the matrix inversion in the characterization above. Not surprisingly, we see the same convergence-complexity trade-off between GD and NM. While NM has a better convergence rate compared to GD, it suffers from a higher per-iteration computational cost. GD, with a proper choice of step-size, converges linearly [12, Thm. 3.12] with $O(\ell)$ per-iteration complexity (disregarding the complexity of gradient oracle). On the other hand, NM, with a proper choice of step-size, has a local quadratic convergence rate [12, Thm. 5.3] with $O(\ell^{3})$ per-iteration complexity, assuming direct inversion (and disregarding the complexity of gradient and Hessian oracles).
@@
where $R=(r_{1},\ldots,r_{j}),\ B=(b_{1},\ldots,b_{j})\in\mathbb{R}^{\ell\times j}$. Hence, $\widetilde{H}_{k}^{-1}$ can be easily computed based on $H_{\mathrm{prior}}^{-1}$ using the Woodbury formula. Different choices of the prior and the linear constraints in the generic approximation scheme above lead to different QNMs. For example, by choosing the so-called secant conditions with $r_{i}=x_{k-i+1}-x_{k-i}$ and $b_{i}=\nabla f(x_{k-i+1})-\nabla f(x_{k-i})$ as linear constraints and $H_{\mathrm{prior}}=I$ as the prior, we derive Anderson mixing with memory $j$ [2], while by using a single secant condition with $j=1$ and choosing $H_{\mathrm{prior}}=\widetilde{H}_{k-1}$ as the prior, we derive QNM with Broyden approximation [11].
In what follows, we use a similar idea and propose the quasi-policy iteration (QPI) algorithm by incorporating a computationally efficient approximation of the “Hessian” $H=I-\gamma P$ in the PI algorithm. We note that the authors in [19, 56, 48] also propose the combination of Anderson mixing with optimal control algorithms. However, the QPI algorithm is fundamentally different in the sense that it approximates the transition matrix $P$ using a different set of constraints that are specific to the optimal control algorithms.
@@ 3.2. QPI Algorithm
The update rule (8) using the approximation (12) is, however, not necessarily a contraction. The same problem also arises in similar algorithms such as Anderson accelerated VI [56] and Nesterov accelerated VI [22]. Here, we follow the standard solution for this problem, that is, safeguarding the QPI update against the standard VI update based on the Bellman error
@@ 3.2.1. The prior
Another interesting choice for the prior in (12) is $P_{\mathrm{prior}}=\widetilde{P}_{k-1}$, i.e., the previous approximation. This leads to a recursive scheme for approximating the transition matrix similar to QNM with Broyden approximation [11]. This can be achieved by choosing an initialization $\widetilde{P}_{-1}$ such that $\widetilde{P}_{-1}\boldsymbol{1}=\boldsymbol{1}$ and defining $\widetilde{G}_{-1}\coloneqq(1-\gamma\widetilde{P}_{-1})^{-1}$ (e.g., $\widetilde{P}_{-1}=\frac{1}{n}E$ and $\widetilde{G}_{-1}=I+\frac{\gamma}{n(1-\gamma)}E$). We note that for this recursive scheme, one might also need to introduce some form of projection to avoid unbounded growth of $\widetilde{P}_{k}$ and $\widetilde{G}_{k}$. In particular, using the fact that $\widetilde{P}_{k}$ is in principle an approximation of a row stochastic matrix, one can use a simple normalization to ensure $\|\widetilde{P}_{k}\|_{\infty}\leq 1$ and $\|\widetilde{G}_{k}\|_{\infty}\leq(1-\gamma)^{-1}$ for all $k$.
@@ 3.3. Extension to model-free control: QPL algorithm
where $\alpha_{k}$ is the diminishing learning rate of the algorithm, e.g., $\alpha_{k}=1/(k+1)$. However, similar to the model-based case, the update rule (20) is not convergent. To address this issue, we consider a convex combination of the update rule (20) with the “standard” synchronous Q-learning (QL) update rule [^3] This is the so-called synchronous update of the Q-function in all state-action pairs in each iteration, corresponding to the parallel sampling model introduced by [29]. [53, 29]
@@ 4. Technical Proofs
In what follows, we show that the proposed algorithm is an instance of semismooth QNM with a structured least-change secant update and a standard backtracking scheme. To this end, observe that we are looking for the unique root $v^{\star}$ of the map g(v)=vT(v)g(v)=v-T(v). First, note that $g$ is piece-wise affine, Lipschitz-continuous, and, in particular, strongly semismooth [17, Prop. 7.4.7]. Moreover, the any generalized Jacobian $H_{k}\in\partial g(v)\subset\{I-\gamma P\;:\;P\geq 0,\ P\boldsymbol{1}=\boldsymbol{1}\}$ is non-singular at any point $v$ (note that for a row stochastic matrix $P$ and $\gamma\in(0,1)$, we have $(I-\gamma P)^{-1}=\sum_{t=0}^{\infty}\gamma^{t}P^{t}$).
@@
besides the structural linear constraints that the true (generalized) Jacobian must satisfy. This implies that the proposed algorithm indeed involves a least-change in $\|\widetilde{H}_{k}-\widetilde{H}_{k-1}\|_{F}$ subject to secant and structural constraints. Now, since $g$ is strongly semismooth, there exists $H_{k}\in\partial g(v_{k})$ such that [17, Def. 7.4.2]
@@
That is, Dennis-Moré condition is satisfied and, by [47, Thm. 4.2], the convergence is locally superlinear.
@@
where $\bar{T}$ is the corresponding true Bellman operator for Q-functions, i.e., $\bar{T}(q)=\mathds{E}[\widehat{T}(q)]$ for each $q$. Then, the convergence of $q_{k}$ to the unique equilibrium $q^{\star}$ of the map $F$ follows from [39, Cor. 3.2]. In particular, observe that $d_{k}\rightarrow 0$ since $\beta_{k}\rightarrow 0$ and $\left\|\Pi_{M}(p_{k})\right\|_{\infty}\leq M$.
@@ 5.1. Model-based algorithms
For model-based algorithms we consider two MDPs: a randomly generated Garnet MDP [3] and the Healthcare MDP [22] with an absorbing state. The proposed QPI algorithm (15) is compared with the following algorithms: VI (value iteration); NVI (VI with Nesterov acceleration) [22]; AVI (VI with Anderson acceleration) [19]; and, PI (policy iteration). For AVI, we use a memory of one leading to a rank-one update (of the identity matrix) for approximating the Hessian so that it is comparable with the rank-one update of the uniform distribution for approximating the transition matrix in QPI. See Appendix B for the exact update rules of NVI and AVI. We note that since NVI and AVI are not guaranteed to converge, we safeguard them using VI (using the same safeguarding rule (13) used for QPI). All the algorithms are initialized by $v_{0}=\boldsymbol{0}$ with termination condition $\left\|v_{k}-T(v_{k})\right\|_{\infty}\leq 10^{-6}$. The results of the simulations are provided in Figures 1 and 2.
@@ 5.2. Model-free algorithms
For model-free algorithms we also consider two MDPs: again a randomly generated Garnet MDP [3] and the Graph MDP [15]. The proposed QPL algorithm (23) with $\alpha_{k}=\frac{1}{(k+1)}$ and $\beta_{k}=\frac{1}{(k+1)^{0.1}}$ is compared with the following algorithms: QL (Q-learning) as in (21) with $\alpha_{k}=\frac{1}{(k+1)}$; SQL (speedy QL) [20]; and, ZQL (zap QL) [15]. See Appendix B for the exact update rules of SQL and ZQL. All the algorithms are initialized by $q_{0}=\boldsymbol{0}_{nm}$ and terminated after $K=10^{4}$ iterations with a synchronous sampling of all state-action pairs at each iteration. For each algorithm, we report the average of the Bellman error $\left\|q_{k}-T(q_{k})\right\|_{\infty}$ over 20 runs of the algorithm. The results of the simulations are provided in Figure 3 and Table 1.

### v4

@@ Abstract.
The authors are with (a) Delft University of Technology, The Netherlands, and (b) University of Toronto, Canada.
This work was partially supported by the European Research Council (ERC) under the grant TRUST-949796, and the NSERC Discovery grant RGPIN-2025-06544.
@@ 1. Introduction
The problem of control, or the decision-making problem as it is also known within the operations research community, has been the subject of much research since the introduction of the Bellman principle of optimality in the late 1950s [bellman1957markovian]. In particular, the connection between control algorithms for Markov decision processes (MDPs) and optimization algorithms has been noticed since the late 1970s [puterman1979convergence]: Value iteration (VI) [bellman1957markovian] can be seen as an instance of gradient descent (GD) algorithm, and policy iteration (PI) [howard1960dynamic] is an instance of the Newton method (NM).
@@
More recent works have used the relationship mentioned above to develop new control algorithms, with faster convergence and/or lower complexity, inspired by their counterparts for solving optimization problems [grand2021convex, vieillard2019connections]. In case of model-based (a.k.a. planning, with access to the model of the MDP) algorithms, the combination of the VI algorithm with Polyak momentum [polyak1964some], Nesterov acceleration [nesterov1983method], and Anderson acceleration [anderson1965iterative] have been explored in [goyal2019first] and [zhang2020globally]. More recently, Halpern anchoring acceleration [halpern1967fixed] has been used to introduce the Anchored VI algorithm [lee2024accelerating], which in particular exhibits an $\operatorname{\mathcal{O}}(1/k)$-rate for large values of discount factor and even for $\gamma=1$. Also of notice is the Generalized Second-Order VI algorithm [kamanchi2021generalized] which applies NM on a smoothed version of the Bellman operator. For model-free (a.k.a. learning, with access to samples from the MDP) algorithms, Speedy Q-Learning [ghavamzadeh2011speedy], Momentum Q-Learning [weng2020momentum], and Nesterov Stochastic Approximation [devraj2019matrix] are among the algorithms that use the idea of momentum for accelerating stochastic GD algorithm [yang2016unified, kidambi2018insufficiency, liu2018accelerating, allen2017katyusha] in order to achieve a better rate of convergence compared to standard Q-learning (QL). Moreover, the Zap Q-Learning algorithm [devraj2017zap] can be thought of as a second-order learning algorithm which was inspired by the stochastic Newton-Raphson (SNR) algorithm [ruppert1985newton].
@@
We note that all of the aforementioned connections have been focused on finite state-action MDPs. When it comes to infinite (continuous) state-action spaces, except in special cases such as linear–quadratic regulators (LQR), one needs to resort to finite-dimensional approximation techniques for computational purposes. This approximation may be at the modeling level by aggregation (discretization) of the state and action spaces, which readily falls into the finite state-action MDP setting [ref:Bert_disc_75, ref:Powell_07]. Alternatively, one may directly approximate the value function via finite parametrization and minimizing (a proxy of) the residual of its fixed-point characterization based on the Bellman principle of optimality [ref:Bert_abstract, szepesvari2022algorithms]. Examples of such include linear parameterization [ref:Bert_temporal, Tsitsiklis97], or nonlinear parameterization with, for instance, neural network architectures [ref:Bert_neuro-dynamic, SCHMIDHUBER2015, sutton2018reinforcement] or max-plus approximation [Bach20, GONCALVES2021109623, ref:FDP_TAC, ref:FDP_NeurIPS, McEn06, amin2025fitted]. We also note that there is an alternative characterization of the original function as the solution to an infinite-dimensional linear program [ref:HL1], paving the way for approximation techniques via finite tractable convex optimization [ref:VanRoy_MP, ref:HL2, ref:Moh_SIOPT]. With this view of the literature, it is worth noting that one can cast almost all of these approximation techniques as the solution to a finite-dimensional fixed-point or convex optimization problem.
@@ 3. Quasi-Policy Iteration (QPI)
where $f:\mathbb{R}^{\ell}\rightarrow\mathbb{R}$ is twice continuously differentiable and strongly convex, and let $g(x)=\nabla f(x)$ and $H(x)=\nabla^{2}f(x)$ be the gradient and Hessian of $f$ evaluated at $x$. Then, by defining $g(v)\coloneqq v-T(v)$, one can see that the VI update rule [bellman1957markovian]
@@
where $\alpha_{k}$ is the step-size. Similarly, by defining $H(v)\coloneqq I-\gamma P^{\pi_{v}}$, the PI update rule [puterman2014markov, Prop. 6.5.1]
@@
The VI algorithm converges linearly with rate $\gamma$ owing to the fact that the operator $T$ is a $\gamma$-contraction in the $\infty$-norm [puterman2014markov, Prop. 6.2.4]. The PI algorithm, on the other hand, outputs the optimal policy in a finite number of iterations [puterman2014markov, Thm. 6.4.2]. Moreover, the algorithm has a local quadratic rate of convergence when initiated in a small enough neighborhood around the optimal solution [bertsekas2022lessons, gargiani2022dynamic]. The faster convergence of PI compared to VI, however, comes with a higher per-iteration computational complexity: The per-iteration complexities of VI and PI are $\operatorname{\mathcal{O}}(n^{2}m)$ and $\operatorname{\mathcal{O}}(n^{2}m+n^{3})$, respectively. The extra $\operatorname{\mathcal{O}}(n^{3})$ complexity is due to the policy evaluation step, i.e., solving a linear system of equations; see also the matrix inversion in the characterization above. Not surprisingly, we see the same convergence-complexity trade-off between GD and NM. While NM has a better convergence rate compared to GD, it suffers from a higher per-iteration computational cost. GD, with a proper choice of step-size, converges linearly [bubeck2015convex, Thm. 3.12] with $O(\ell)$ per-iteration complexity (disregarding the complexity of gradient oracle). On the other hand, NM, with a proper choice of step-size, has a local quadratic convergence rate [bubeck2015convex, Thm. 5.3] with $O(\ell^{3})$ per-iteration complexity, assuming direct inversion (and disregarding the complexity of gradient and Hessian oracles).
@@
where $R=(r_{1},\ldots,r_{j}),\ B=(b_{1},\ldots,b_{j})\in\mathbb{R}^{\ell\times j}$. Hence, $\widetilde{H}_{k}^{-1}$ can be easily computed based on $H_{\mathrm{prior}}^{-1}$ using the Woodbury formula. Different choices of the prior and the linear constraints in the generic approximation scheme above lead to different QNMs. For example, by choosing the so-called secant conditions with $r_{i}=x_{k-i+1}-x_{k-i}$ and $b_{i}=\nabla f(x_{k-i+1})-\nabla f(x_{k-i})$ as linear constraints and $H_{\mathrm{prior}}=I$ as the prior, we derive Anderson mixing with memory $j$ [anderson1965iterative], while by using a single secant condition with $j=1$ and choosing $H_{\mathrm{prior}}=\widetilde{H}_{k-1}$ as the prior, we derive QNM with Broyden approximation [broyden1965class].
In what follows, we use a similar idea and propose the quasi-policy iteration (QPI) algorithm by incorporating a computationally efficient approximation of the “Hessian” $H=I-\gamma P$ in the PI algorithm. We note that the authors in [geist2018anderson, zhang2020globally, sun2021damped] also propose the combination of Anderson mixing with optimal control algorithms. However, the QPI algorithm is fundamentally different in the sense that it approximates the transition matrix $P$ using a different set of constraints that are specific to the optimal control algorithms.
@@ 3.2. QPI Algorithm
The update rule (8) using the approximation (12) is, however, not necessarily a contraction. The same problem also arises in similar algorithms such as Anderson accelerated VI [zhang2020globally] and Nesterov accelerated VI [goyal2019first]. Here, we follow the standard solution for this problem, that is, safeguarding the QPI update against the standard VI update based on the Bellman error
@@ 3.2.1. The prior
Another interesting choice for the prior in (12) is $P_{\mathrm{prior}}=\widetilde{P}_{k-1}$, i.e., the previous approximation. This leads to a recursive scheme for approximating the transition matrix similar to QNM with Broyden approximation [broyden1965class]. This can be achieved by choosing an initialization $\widetilde{P}_{-1}$ such that $\widetilde{P}_{-1}\boldsymbol{1}=\boldsymbol{1}$ and defining $\widetilde{G}_{-1}\coloneqq(1-\gamma\widetilde{P}_{-1})^{-1}$ (e.g., $\widetilde{P}_{-1}=\frac{1}{n}E$ and $\widetilde{G}_{-1}=I+\frac{\gamma}{n(1-\gamma)}E$). We note that for this recursive scheme, one might also need to introduce some form of projection to avoid unbounded growth of $\widetilde{P}_{k}$ and $\widetilde{G}_{k}$. In particular, using the fact that $\widetilde{P}_{k}$ is in principle an approximation of a row stochastic matrix, one can use a simple normalization to ensure $\|\widetilde{P}_{k}\|_{\infty}\leq 1$ and $\|\widetilde{G}_{k}\|_{\infty}\leq(1-\gamma)^{-1}$ for all $k$.
@@ 3.3. Extension to model-free control: QPL algorithm
where $\alpha_{k}$ is the diminishing learning rate of the algorithm, e.g., $\alpha_{k}=1/(k+1)$. However, similar to the model-based case, the update rule (20) is not convergent. To address this issue, we consider a convex combination of the update rule (20) with the “standard” synchronous Q-learning (QL) update rule [^3] This is the so-called synchronous update of the Q-function in all state-action pairs in each iteration, corresponding to the parallel sampling model introduced by [kearns1998finite]. [watkins1992q, kearns1998finite]
@@ 4. Technical Proofs
In what follows, we show that the proposed algorithm is an instance of semismooth QNM with a structured least-change secant update and a standard backtracking scheme. To this end, observe that we are looking for the unique root $v^{\star}$ of the map g(v)=vT(v)g(v)=v-T(v). First, note that $g$ is piece-wise affine, Lipschitz-continuous, and, in particular, strongly semismooth [facchinei2003finite, Prop. 7.4.7]. Moreover, the any generalized Jacobian $H_{k}\in\partial g(v)\subset\{I-\gamma P\;:\;P\geq 0,\ P\boldsymbol{1}=\boldsymbol{1}\}$ is non-singular at any point $v$ (note that for a row stochastic matrix $P$ and $\gamma\in(0,1)$, we have $(I-\gamma P)^{-1}=\sum_{t=0}^{\infty}\gamma^{t}P^{t}$).
@@
besides the structural linear constraints that the true (generalized) Jacobian must satisfy. This implies that the proposed algorithm indeed involves a least-change in $\|\widetilde{H}_{k}-\widetilde{H}_{k-1}\|_{F}$ subject to secant and structural constraints. Now, since $g$ is strongly semismooth, there exists $H_{k}\in\partial g(v_{k})$ such that [facchinei2003finite, Def. 7.4.2]
@@
That is, Dennis-Moré condition is satisfied and, by [sun1997newton, Thm. 4.2], the convergence is locally superlinear.
@@
where $\bar{T}$ is the corresponding true Bellman operator for Q-functions, i.e., $\bar{T}(q)=\mathds{E}[\widehat{T}(q)]$ for each $q$. Then, the convergence of $q_{k}$ to the unique equilibrium $q^{\star}$ of the map $F$ follows from [perkins2013asynchronous, Cor. 3.2]. In particular, observe that $d_{k}\rightarrow 0$ since $\beta_{k}\rightarrow 0$ and $\left\|\Pi_{M}(p_{k})\right\|_{\infty}\leq M$.
@@ 5.1. Model-based algorithms
For model-based algorithms we consider two MDPs: a randomly generated Garnet MDP [archibald1995generation] and the Healthcare MDP [goyal2019first] with an absorbing state. The proposed QPI algorithm (15) is compared with the following algorithms: VI (value iteration); NVI (VI with Nesterov acceleration) [goyal2019first]; AVI (VI with Anderson acceleration) [geist2018anderson]; and, PI (policy iteration). For AVI, we use a memory of one leading to a rank-one update (of the identity matrix) for approximating the Hessian so that it is comparable with the rank-one update of the uniform distribution for approximating the transition matrix in QPI. See Appendix B for the exact update rules of NVI and AVI. We note that since NVI and AVI are not guaranteed to converge, we safeguard them using VI (using the same safeguarding rule (13) used for QPI). All the algorithms are initialized by $v_{0}=\boldsymbol{0}$ with termination condition $\left\|v_{k}-T(v_{k})\right\|_{\infty}\leq 10^{-6}$. The results of the simulations are provided in Figures 1 and 2.
@@ 5.2. Model-free algorithms
For model-free algorithms we also consider two MDPs: again a randomly generated Garnet MDP [archibald1995generation] and the Graph MDP [devraj2017zap]. The proposed QPL algorithm (23) with $\alpha_{k}=\frac{1}{(k+1)}$ and $\beta_{k}=\frac{1}{(k+1)^{0.1}}$ is compared with the following algorithms: QL (Q-learning) as in (21) with $\alpha_{k}=\frac{1}{(k+1)}$; SQL (speedy QL) [ghavamzadeh2011speedy]; and, ZQL (zap QL) [devraj2017zap]. See Appendix B for the exact update rules of SQL and ZQL. All the algorithms are initialized by $q_{0}=\boldsymbol{0}_{nm}$ and terminated after $K=10^{4}$ iterations with a synchronous sampling of all state-action pairs at each iteration. For each algorithm, we report the average of the Bellman error $\left\|q_{k}-T(q_{k})\right\|_{\infty}$ over 20 runs of the algorithm. The results of the simulations are provided in Figure 3 and Table 1.